# Retrieval Augmented Generation or Long-Context LLMs? A Comprehensive Study and Hybrid Approach

### v1

@@ 2 Related Work
Retrieval-augmented generation. Augmenting LLMs with relevant information retrieved from various sources Lewis et al. (2020), i.e., RAG, has been successful in complementing LLMs with external knowledge. RAG achieves good performance on various of tasks like language modeling Khandelwal et al. (2019); Shi et al. (2023) and QA Guu et al. (2020); Izacard and Grave (2020), with a significantly lower computation cost Borgeaud et al. (2022). Related to but different from our work, recently works augment RAG with correction Yan et al. (2024), critique Asai et al. (2023), or verification Li et al. (2023) to improve retrieval quality on knowledge-intensive tasks.
@@ 5.4 Results on synthetic data
This “PassKey” dataset presents a needle-in-a-haystack test, where a sentence with a passkey (e.g. ‘‘the passkey is 123456’’) is hidden within chunks of irrelevant text, and the model is asked to answer the question ‘‘What is the passkey’’. The task requires strong retrieval capability. On this dataset, RAG achieves 80.34% accuracy, outperforming LC, which gets 65.25% using Gemini-1.5-Pro. However, if the query is slightly modified as ‘‘What is the special token hidden inside the texts’’, RAG accuracy sharply drops to only 4.58%, while LC keeps roughly the same (69.32%). Another example: if the chunks contain two passkeys and the query is ‘‘Which passkey is larger? First or second?’’, then RAG (47.63%) under-performs LC (64.24%) as well. These examples demonstrate that the evaluation results highly subjects to artifacts in dataset construction, showing limitation of synthetic testing.

### v2

@@ 2 Related Work
Retrieval-augmented generation. Augmenting LLMs with relevant information retrieved from various sources Lewis et al. (2020) has been successful in complementing LLMs with external knowledge. RAG achieves good performance on tasks like language modeling Khandelwal et al. (2019); Shi et al. (2023) and QA Guu et al. (2020); Izacard and Grave (2020), with a significantly lower computation cost Borgeaud et al. (2022). Related to but different from our work, recently works augment RAG with correction Yan et al. (2024), critique Asai et al. (2023), verification Li et al. (2023), or adaptive search Wang et al. (2023); Cheng et al. (2024); Jeong et al. (2024) to improve retrieval quality on knowledge-intensive tasks.
@@ 5.4 Results on synthetic data
This “PassKey” dataset presents a needle-in-a-haystack test, where a sentence with a passkey (e.g. ‘‘the passkey is 123456’’) is hidden within chunks of irrelevant text, and the model is asked to answer the question ‘‘What is the passkey’’. The task requires strong retrieval capability. On this dataset, RAG achieves 80.34% accuracy, outperforming LC, which gets 65.25% using Gemini-1.5-Pro. However, if the query is slightly modified as ‘‘What is the special token hidden inside the texts’’, RAG accuracy sharply drops to only 4.58%, while LC keeps roughly the same (69.32%). Another example: if the chunks contain two passkeys and the query is ‘‘Which passkey is larger? First or second?’’, then RAG (47.63%) under-performs LC (64.24%) as well. Tab. 3 summarizes the results, which demonstrates that the evaluation highly subjects to artifacts in dataset construction, showing limitation of synthetic testing.
RAG LC
Original 80.34 65.25
Variant-1: “special token” 4.58 69.32
Variant-2: “which is larger” 47.63 64.24
Table 3: Synthetic dataset may unconsciously contain artifacts that influence the comparison results.
5.5 Exclusion of LLM’s internal knowledge
Ideally, the comparison in this paper should exclude the model’s internal knowledge (i.e., parametric knowledge) so that the model’s performance are solely based on its capability to understand long contexts. In our study, this internal knowledge is excluded by utilizing the prompt “based only on the provided passage”, which we empirically find is a simple yet effective method. Here we discuss the effectiveness of this method, as well as alternative methods to exclude external knowledge.
First, we validate the effectiveness of the simple prompt “based only on the provided passage”. Tab. 4 compares the performance (long-context) of Gemini-1.5-Pro with and without this prompt. As shown, using this prompt consistently limits the model’s performance (average performance drops from 50.57 to 45.53), which indicates that using this simple instruction can already effectively limit the usage of the model’s parametric knowledge.
without
"based only on …"
with
"based only on …"
NarrativeQA 36.35 32.76
Qasper 50.69 47.83
MultiFieldQA 56.07 52.33
HotpotQA 66.47 61.85
2WikiMQA 68.97 62.96
Musique 54.56 40.22
QMSum 20.87 20.73
En.QA 49.20 43.08
En.MC 90.83 85.57
Avg 50.57 45.53
Table 4: Comparison of the long-context performance of Gemini-1.5-Pro, using the prompt with and without “based only on the provided passage”.
Second, as an alternative method to exclude internal knowledge, we remove the questions where the model can correctly answer without any contexts (i.e., commonsense questions), and report the model’s performance only on the non-commonsense questions. Tab. 5 shows the performance of Gemini-1.5-Pro and GPT-3.5-Turbo on all the questions from the MuSiQue dataset, as well as their performance on the non-commonsense subset [^3] Different models may learn different internal knowledge, resulting in different numbers of non-commonsense questions. For example, GPT-3.5-Turbo gets 14.53 performance on MuSiQue while Gemini-1.5-Pro gets 23.58 using only internal knowledge,. As shown, after excluding the commonsense questions, the trend remains the same.
all questions w/o commonsense
Gemini GPT-3.5 Gemini GPT-3.5
# questions 200 200 133 150
LC 40.22 17.92 31.76 13.00
RAG 26.56 16.41 15.51 13.05
Self-Route 40.66 34.44 31.32 19.76
answerable % 58.50 47.00 52.63 45.33
token % 56.14 65.08 48.46 53.43
Table 5: Results on MuSiQue on all questions, and on the subset of non-commonsense questions (i.e., excluding questions that can be answered without contexts).
That said, a more thorough study to explore various methods for controlling the usage of model’s internal knowledge, and to study the source of internal knowledge (e.g. LLM"s world knowledge or dataset leakage), will be valuable future work, which we hope can be further investigated.
@@ 5.5 Exclusion of LLM’s internal knowledge
Acknowledgements
We would like to thank Weize Kong, Tao Chen, Jeffrey Dudek and Spurthi Amba Hombaiah for their helpful comments and suggestions, as well as the anonymous reviewers for the valuable discussions.